Systems Staff Meeting Notes 07/22/02

Systems Staff Meeting Notes
July 22, 2002

Present: Rich Bennett, Cynthia Bowker, Bill Covey, Suzy Covey, Debra Harris, Bobbie Meng, Hilton Soto, Ying Tang
  1. Bobbie got caught in a vicious hailstorm in Jacksonville over the weekend.
  2. China has been experiencing some strange weather this summer.
  3. There are plagues of locusts in Africa.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Ordering situation has become complex – in general, routine maintenance items will be ordered as usual. New items may take up to a week before they are ordered. All new orders now have to be approved by the directors. We will process all orders as usual – no holding them until time to be approved. How long it takes after we process the orders is unknown.
  2. Ariel – PDF problems – have not heard anything from ILL.
  1. MMM – had no problem with the liaisons setting a quick lockout on staff screen savers. Will put the word out this week.
  1. Electrical power on campus is currently being revamped. It will continue to be unstable throughout the summer. For this reason, it’s important to keep up with your personal backups.
  2. Server backups – media, drive and operator problems – the weekly backups appear to be using more tape. Last week it took until Wednesday to get the Friday backups done. Friday and the NT12 backups are still running today. Bobbie put in a new set of tapes. Something is odd – the byte count is high – something has changed. Eventually, everything will work.
  3. Accounts locked by outside probes – filters are now going to be put on the campus Internet POP. All of the attacks that hit us were from outside of UF. This filter would eliminate such attacks 100%. They wanted to install the filter two weeks after fall classes start, but Bill wanted it to be done before the fall term – in case we run into any difficulties. Right now it’s scheduled to be implemented on August 18. This will eliminate network neighborhood browsing and network administration from outside campus unless it comes through VPN tunnels. VPN requires new software to be installed. This will have an influence on Suzy’s work from home.
  4. DMZ is being created on campus – they are easing it in on the firewall. It’ll be set up so our things stay the same, but we’ll be in front of one firewall and behind another. They are not ready to announce this move yet – it has not happened as of yet.
  5. Bill is pushing to get all user folders on NT2R.
Software Projects Status
  1. Cynthia is working on the SW/HW online request form.
  2. Ying and Winston had a review with Erich – will use a button instead of email notification in the tracking database.
  3. DLC requested a change in the CD Lister program.
  4. EShare Net Agent – Cynthia worked a little bit on the transcript search – she will talk to Bill to discuss further.
  5. Bill is requesting more Trouble Ticket and IRDB mods.
DP Coordinator
  1. Dan Reboussin is requesting the use of a laptop in the West conference room. Need to check to see if there is one already assigned to that room. Seems like Gary Cornwell has one that they use for that purpose, but he is away at camp and has the laptop with him. Will investigate.

Back to the Systems Homepage

Last updated July 29, 2002
by Debra Harris